OpenStudy (anonymous):

HELP I"LL MEDAL!!!!!! Do the expressions in parts A and B have a common factor? If so, what is the common factor? A x2 + 2x - 24 B x2 – 8x + 16

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

no

OpenStudy (skullpatrol):

Try the factoring pattern a^2 - 2ab + b^2 for x^2 - 8x + 16.

OpenStudy (anonymous):

no I mean I think its no

OpenStudy (skullpatrol):

Recall a^2 - 2ab + b^2 = ( a - b )^2

OpenStudy (triciaal):

part A what are the factors of -24 that add up to +2 pay attention to the signs

OpenStudy (triciaal):

can you rewrite group and factor ?

OpenStudy (triciaal):

part B can you factorize?

OpenStudy (skullpatrol):

So x^2 - 8x + 16 = x^2 - 2*4x + 4^2